# LQG inspired spacetimes as solutions of the Einstein equations

### v1

@@ I Introduction
The theory of general relativity (GR) proposed by Albert Einstein in 1915 represents a revolution in modern thinking regarding the concept of gravity, which is no longer interpreted as a force and is now interpreted as a geometric effect reactive to matter fields. This is one of the most successful physical theories in terms of experimental and observational verifications [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Therefore, its acceptance as a starting point for the investigation of gravitational systems is unanimous among experts in the field. Several spherically symmetric models presenting geodesic singularity problems were proposed within this important gravitational theory, describing the so-called black hole (BH) solutions [11]. In this context, the first and simplest proposed solution was that of Schwarzschild, who described a vacuum solution for a spherically symmetric and static BH. Subsequently, several other singular models emerged for spherically symmetric and static systems and spherically symmetric and stationary systems. Although they commonly exhibit singularities in their interior, the main characteristic of a BH is the presence of an event horizon, a no-return surface that causally separates distinct regions of spacetime.
@@
In addition to these systems that presented problems regarding the presence of singularity, another class of models, now without the presence of singularity, was initially proposed by James Bardeen around 1968 to represent the geometry of a regular black hole (RBH) [12]. Later, around 2000, Beato and Garcia obtained the source of matter for this prototype, which was nonlinear electrodynamics (NED), in which the regularization parameter begins to be interpreted as the charge of the magnetic monopole [13]. Still in the context of these RBH solutions, several other models were proposed, such as the electrically charged case analyzed by Rodrigues and Silva [14] among others [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
@@
Still in the scenario of regular solutions, recently Simpson and Visser proposed a new class of compact objects that describe some variations of wormhole (WH) solutions depending on the relationship between the model parameters, which are the mass m and the radius of the throat a [27]. Such objects were called black bounces (BB) and can present some possible configurations according to the relationship between the model parameters. Thus, if $a=2m$ we have the formation of a unidirectional WH, if $a>2m$ we have the case of a bidirectional WH, and for $a<2m$ we have the formation of a WH throat hidden by two symmetric event horizons. The source of matter associated with this regularized gravitational system that does not satisfy vacuum conditions was later found to be a combination of NED and phantom scalar fields [28, 29, 30]. Thus, after this class of BB solutions emerged, several other gravitational scenarios were explored to verify possible new BB solutions. Scenarios include k-essence theories [31, 32, 33], modified gravity [34, 35, 36, 37], conformal gravity [38], teleparallel gravity [39], black strings [40, 41, 42, 43], spherically symmetric and stationary backgrounds [44, 45, 46], among others [47, 48, 49, 50, 51]. These BB models are being analyzed through theoretical and observational tests such as for gravitational lensing effects in the weak and strong field regime [52, 53, 54, 55, 56, 57, 58, 59, 60, 61], quasi-normal modes, shadows, and optical appearance and scattering [62, 63, 64, 65, 66].
@@
From this perspective, Loop Quantum Gravity (LQG) emerges as a compelling framework for describing gravitational systems beyond GR, particularly in regimes where quantum effects become significant. It provides tools to investigate compact objects, both with and without geodesic singularities, providing insight into the quantum structure of spacetime. A key feature is the polymer representation [67, 68], which enables the mapping of singular spacetimes to fully regular geometries through effective quantum corrections.
@@
Although LQG does not yet constitute a complete quantum theory of spacetime, particularly near classical singularities, its applications have been predominantly concentrated in semiclassical regimes. Substantial progress has been made in modeling singularity resolution and building effective quantum-corrected geometries [69, 70, 71, 72, 73, 74], including configurations involving WH structures [75, 76]. In this context, [77] proposed a procedure to introduce LQG-inspired corrections into already regular spacetimes. This method, based on a coordinate transformation reminiscent of the approach by Simpson and Visser [27], yields regular geometries that incorporate nontrivial quantum gravitational effects, as exemplified in [78]. Notably, in the latter reference, the resulting RBHs can be interpreted as structures supported by sources modeled through NED within the framework of GR. These developments broaden the applicability of LQG-based models, providing new pathways for constructing physically realistic, non-singular solutions within a semiclassical gravitational framework.
Thus, we consider the Simpson-Visser spacetime as a starting point to test the formalism required to find the source fields of BHs with LQG corrections as solutions of GR. We constructed sources for electrically and magnetically charged solutions, which were a combination of NED and phantom scalar fields, and then the energy conditions were constructed and analyzed.
@@
We adopt the metric signature (+,,,)(+,-,-,-)( + , - , - , - ). We shall work in geometrodynamics units where $18\pi G=\hbar=c=$ 1.
@@ II Field source to a general spacetime
This type of problem also arises in the case of BB spacetimes, such as the Simpson-Visser model. As is usually done in the Simpson-Visser model, we will check whether the coupling of gravitational theory together with NED and a phantom scalar field is capable of generating this spacetime.
@@ II.1 Magnetic case
$\frac{-r^{3}g\left(rf^{\prime}g^{\prime}+2f\left(rg^{\prime\prime}+g^{\prime}\right)\right)-2r^{3}g^{2}f^{\prime}+r^{4}fg^{\prime 2}}{4r^{4}g^{2}}=\frac{L}{2}-\frac{2q^{2}L_{F}}{r^{4}}+V+\epsilon f\phi^{\prime 2},$ (15)
@@
$L_{F}(r)$ $=$ $\frac{r^{2}\left(rg\left(rf^{\prime}g^{\prime}-2f\left(g^{\prime}-rg^{\prime\prime}\right)\right)+2g^{2}\left(rf^{\prime}+f\left(4r^{2}\epsilon\phi^{\prime 2}-2\right)+2\right)-r^{2}fg^{\prime 2}\right)}{8q^{2}g^{2}}.$ (18)
@@
$\phi^{\prime}(r)$ $=$ $\frac{i\sqrt{gf^{\prime}-fg^{\prime}}}{\sqrt{2r\epsilon fg}}.$ (20)
@@ II.2 Electric case
$F^{10}=\frac{q}{L_{F}r^{2}}\sqrt{\frac{f}{g}},\quad\mbox{and}\quad F=-\frac{2q^{2}}{r^{4}L_{F}^{2}}.$ (21)
@@
$-\frac{f^{\prime}}{r}-\frac{f}{r^{2}}+\frac{1}{r^{2}}=\frac{L}{2}+\frac{2q^{2}}{r^{4}L_{F}}+\epsilon f\phi^{\prime 2}+V,$ (22)
$-\frac{f\left(\frac{rg^{\prime}}{g}+1\right)}{r^{2}}+\frac{1}{r^{2}}=\frac{L}{2}+\frac{2q^{2}}{r^{4}L_{F}}-\epsilon f\phi^{\prime 2}+V,$ (23)
@@
$L_{F}(r)=\frac{8q^{2}g^{2}}{r^{2}\left(rg\left(rf^{\prime}g^{\prime}-2f\left(g^{\prime}-rg^{\prime\prime}\right)\right)+2g^{2}\left(rf^{\prime}+f\left(-4r^{2}\phi^{\prime 2}-2\right)+2\right)-r^{2}fg^{\prime 2}\right)},$ (26)
@@
Although obtained independently through the Einstein equations, the electromagnetic functions $L$ and $L_{F}$, for both the electric and magnetic cases, are related to each other and must obey the consistency relation:
$L_{F}-\frac{\partial L}{\partial F}=L_{F}-\frac{\partial L}{\partial r}\left(\frac{\partial F}{\partial r}\right)^{-1}=0.$ (27)
@@
$P=P^{\mu\nu}P_{\mu\nu}=L_{F}^{2}F=-\frac{2q^{2}}{r^{4}}.$ (28)
In general, the above function is much easier to invert than the scalar $F$, equation (21).
@@ III Simpson–Visser Solution
In order to test the consistency of the method, we will first consider the Simpson-Visser BB solution, which has a bounce structure similar to LQG-inspired solutions. Usually, the line element that describes this spacetime is written as [27]
@@
Despite presenting itself differently from the equation (1), through a coordinate transformation, we can rewrite this line element as the equation (1) with the functions $f(r)$ and $g(r)$ given by [28]
@@
$L(r)$ $=$ $\frac{12a^{2}m}{5r^{5}},\quad L_{F}(r)=\frac{3a^{2}m}{2q^{2}r},\quad F(r)=\frac{2q^{2}}{r^{4}}.$ (33)
With appropriate adaptations of the coordinates and notation, the results are compatible with those obtained in previous works [28, 29, 30]. The electromagnetic quantities satisfy the relation (27).
@@
[…]
@@
$L(r)=-\frac{18a^{2}m}{5r^{5}},\quad L_{F}(r)=\frac{2q^{2}r}{3a^{2}m},\quad F(r)=-\frac{9a^{4}m^{2}}{2q^{2}r^{6}}.$ (35)
@@
Once again, the result is consistent with those already existing in the literature [79]. This shows us that the method is effective in obtaining the source of some types of BBs in GR.
It is important to emphasize that in some cases it is not possible to represent the source of solutions using a scalar field that is purely canonical or purely phantom, thus requiring the consideration of partially phantom scalar fields [80, 81, 82, 83, 37]. This type of model demands a greater degree of freedom in the scalar field, thereby introducing additional functions to be determined.
@@ IV Holonomy corrected Schwarzschild BH
The Holonomy corrected Schwarzschild BH is described by the line element (1) with [84]
@@
where $0<r_{0}<2m$ represents the radius of the throat of a WH with minimal area $4\pi r_{0}^{2}$ [85].
Despite the metric coefficient being different, the causal structure of the Simpson-Visser solution is very similar to Holonomy corrected Schwarzschild BH. Therefore, it is possible that this spacetime, like the Simpson-Visser solution, can also be found from GR with the appropriate sources.
@@
$L(r)$ $=$ $\frac{r_{0}(9m+4r)}{6r^{4}},\quad L_{F}(r)=\frac{r_{0}(3m+r)}{4q^{2}},\quad F(r)=\frac{2q^{2}}{r^{4}}.$ (39)
@@
$L(F)=\frac{r_{0}\left(4\sqrt[4]{2}F^{3/4}\sqrt{q}+9Fm\right)}{12q^{2}},\qquad V(\phi)=\frac{\cos^{6}\left(\frac{\phi}{\sqrt{2}}\right)\left(3m\cos\left(\sqrt{2}\phi\right)+3m+4r_{0}\right)}{24r_{0}^{3}}.$ (40)
In the limit of weak fields, $F\approx 0$, the dominant term is $F^{3/4}$, which is nonlinear.
@@
$L(r)=-\frac{r_{0}(9m+2r)}{6r^{4}},\quad L_{F}(r)=\frac{4q^{2}}{3mr_{0}+rr_{0}},\quad F(r)=-\frac{r_{0}^{2}(3m+r)^{2}}{8q^{2}r^{4}}.$ (41)
@@
$L(P)=\frac{r_{0}\left(9mP-2\sqrt[4]{2}(-P)^{3/4}\sqrt{q}\right)}{12q^{2}}.$ (42)

### v2

@@ I Introduction
The theory of general relativity (GR) proposed by Albert Einstein in 1915 represents a revolution in modern thinking regarding the concept of gravity, which is no longer interpreted as a force and is now interpreted as a geometric effect reactive to matter fields. This is one of the most successful physical theories in terms of experimental and observational verifications [INTRO1, INTRO2, INTRO3, INTRO4, INTRO5, INTRO6, INTRO7, INTRO8, INTRO9, INTRO10]. Therefore, its acceptance as a starting point for the investigation of gravitational systems is unanimous among experts in the field. Several spherically symmetric models presenting geodesic singularity problems were proposed within this important gravitational theory, describing the so-called black hole (BH) solutions [Schwarzschild:1916uq]. In this context, the first and simplest proposed solution was that of Schwarzschild, who described a vacuum solution for a spherically symmetric and static BH. Subsequently, several other singular models emerged for spherically symmetric and static systems and spherically symmetric and stationary systems. Although they commonly exhibit singularities in their interior, the main characteristic of a BH is the presence of an event horizon, a no-return surface that causally separates distinct regions of spacetime.
@@
In addition to these systems that presented problems regarding the presence of singularity, another class of models, now without the presence of singularity, was initially proposed by James Bardeen around 1968 to represent the geometry of a regular black hole (RBH) [INTRo11]. Later, around 2000, Beato and Garcia obtained the source of matter for this prototype, which was nonlinear electrodynamics (NED), in which the regularization parameter begins to be interpreted as the charge of the magnetic monopole [INTRO12]. Still in the context of these RBH solutions, several other models were proposed, such as the electrically charged case analyzed by Rodrigues and Silva [INTRO13] among others [INTRO14, INTRO15, INTRO16, INTRO17, INTRO18, INTRO19, INTRO20, INTRO21, INTRO22, INTRO23, INTRO24, INTRO25].
@@
Still in the scenario of regular solutions, recently Simpson and Visser proposed a new class of compact objects that describe some variations of wormhole (WH) solutions depending on the relationship between the model parameters, which are the mass m and the radius of the throat a [Simpson:2018tsi]. Such objects were called black bounces (BB) and can present some possible configurations according to the relationship between the model parameters. Thus, if $a=2m$ we have the formation of a unidirectional WH, if $a>2m$ we have the case of a bidirectional WH, and for $a<2m$ we have the formation of a WH throat hidden by two symmetric event horizons. The source of matter associated with this regularized gravitational system that does not satisfy vacuum conditions was later found to be a combination of NED and phantom scalar fields [Canate:2022gpy, Bronnikov:2021uta, Rodrigues:2023vtm]. Thus, after this class of BB solutions emerged, several other gravitational scenarios were explored to verify possible new BB solutions. Scenarios include k-essence theories [CF1, CF2, CF3], modified gravity [GRAVMOD1, GRAVMOD2, GRAVMOD3, GRAVMOD4], conformal gravity [GRAVCONFORM], teleparallel gravity [GRAVTELL], black strings [GEO1, GEO2, GEO3, Geo4], spherically symmetric and stationary backgrounds [ROTA1, ROTA2, GEO2], among others [OUTROS1, OUTROS2, OUTROS3, outros4, outros5]. These BB models are being analyzed through theoretical and observational tests such as for gravitational lensing effects in the weak and strong field regime [BBOBS1, BBOBS2, BBOBS3, BBOBS4, BBOBS5, BBOBS6, BBOBS7, BBOBS8, BBOBS9, BBOBS10, Heidari:2024bkm], quasi-normal modes, shadows, and optical appearance and scattering [BBOBS11, BBOBS12, BBOBS13, BBOBS14, BBOBS15].
@@
From this perspective, Loop Quantum Gravity (LQG) emerges as a compelling framework for describing gravitational systems beyond GR, particularly in regimes where quantum effects become significant. It provides tools to investigate compact objects, both with and without geodesic singularities, providing insight into the quantum structure of spacetime. A key feature is the polymer representation [Modesto:2008im, Laddha:2010wp], which enables the mapping of singular spacetimes to fully regular geometries through effective quantum corrections.
@@
Although LQG does not yet constitute a complete quantum theory of spacetime, particularly near classical singularities, its applications have been predominantly concentrated in semiclassical regimes. Substantial progress has been made in modeling singularity resolution and building effective quantum-corrected geometries [LQG1, LQG2, LQG3, LQG4, LQG5, Lewandowski:2022zce], including configurations involving WH structures [Muniz:2024jzg, deSSilva:2024gdc]. In this context, [LQG6] proposed a procedure to introduce LQG-inspired corrections into already regular spacetimes. This method, based on a coordinate transformation reminiscent of the approach by Simpson and Visser [Simpson:2018tsi], yields regular geometries that incorporate nontrivial quantum gravitational effects, as exemplified in [Muniz:2024wiv]. Notably, in the latter reference, the resulting RBHs can be interpreted as structures supported by sources modeled through NED within the framework of GR. These developments broaden the applicability of LQG-based models, providing new pathways for constructing physically realistic, non-singular solutions within a semiclassical gravitational framework.
The effective matter sector reproducing a given LQG-inspired metric is not unique. This metric itself is already provided as a smooth effective geometry that encodes the coarse-grained influence of the underlying discrete quantum structure of spacetime [BenAchour:2018khr]. Since these geometries describe a continuum regime in which the microscopic LQG degrees of freedom are no longer resolved, and in which a matter sector is required at sufficiently large scales [Kelly:2020uwj], classical GR provides a natural framework in which the metric can be reinterpreted as inducing an effective stress–energy tensor. Our aim is therefore to provide a consistent and minimal classical realization that embeds the LQG-inspired geometry within GR. This reconstruction is useful because it ensures that the spacetime satisfies Einstein’s equations, permits the discussion of physical properties in a familiar macroscopic setting, and yields a consistent framework for perturbative analyses. Although different effective decompositions may exist, as long as they reproduce the same background geometry, the geometric sector of the perturbations – typically the physically relevant one – remains determined by the metric itself. We adopt thus a GR non-vacuum reconstruction (phantom scalar + NED) for the LQG-inspired geometry because it allows a direct perturbative treatment with standard GR tools. The linearized system reduces to second-order wave equations with well-posed boundary conditions, so QNMs, possible echoes, and stability windows can be computed with standard time and frequency domain numerical solvers [Duran-Cabaces:2025sly, Franzin:2023slm]. In contrast, modified gravity models often lead to higher-order or coupled equations, making a well-posed formulation harder and requiring custom numerics.
Thus, we consider the Simpson-Visser spacetime as a starting point to test the formalism required to find the source fields of BHs with LQG corrections as solutions of GR. We constructed sources for electrically and magnetically charged solutions, which were a combination of NED and phantom scalar fields, and then the energy conditions were constructed and analyzed. In our approach, LQG-inspired spacetimes are not treated as fundamentally quantum, but rather as effective geometries that can be consistently described in terms of classical GR. Once such metrics are specified, the corresponding stress-energy tensor can be reconstructed, revealing that the necessary sources may be effectively realized by a phantom scalar field in combination with NED. These classical fields thus provide an effective representation of the energy–momentum content needed to sustain quantum-inspired corrections, as also confirmed by the Simpson-Visser case taken here as a benchmark example. Understanding these sources can be fundamental when studying the properties of the solutions. By interpreting the sources, we can investigate in greater depth the perturbations of the metric and of the source fields. This can reveal more information about the stability of these spacetimes than merely analyzing stability with test fields [Franzin:2023slm].
@@
We adopt the metric signature (+,,,)(+,-,-,-). We shall work in geometrodynamics units where $8\pi G=\hbar=c=1$.
@@ II Field source to a general spacetime
In general, it is always possible to find metrics as solutions of Einstein equations for a given energy density and generic pressures. However, what we seek here are more concrete examples of sources, rather than generic ones, that can generate such solutions. This type of problem also arises in the case of BB spacetimes, such as the Simpson-Visser model. As is usually done in the Simpson-Visser model, we will check whether the coupling of gravitational theory together with NED and a phantom scalar field is capable of generating this spacetime.
@@ II.1 Magnetic case
$\frac{-r^{3}g\left(rf^{\prime}g^{\prime}+2f\left(rg^{\prime\prime}+g^{\prime}\right)\right)-2r^{3}g^{2}f^{\prime}+r^{4}fg^{\prime 2}}{4r^{4}g^{2}}=\frac{L}{2}-\frac{2q_{m}^{2}L_{F}}{r^{4}}+V+\epsilon f\phi^{\prime 2},$ (15)
@@
$L_{F}(r)$ $=$ $\frac{r^{2}\left(rg\left(rf^{\prime}g^{\prime}-2f\left(g^{\prime}-rg^{\prime\prime}\right)\right)+2g^{2}\left(rf^{\prime}+f\left(4r^{2}\epsilon\phi^{\prime 2}-2\right)+2\right)-r^{2}fg^{\prime 2}\right)}{8q_{m}^{2}g^{2}}.$ (18)
@@
$\phi^{\prime}(r)$ $=$ $\frac{i\sqrt{gf^{\prime}-fg^{\prime}}}{\sqrt{2r\epsilon fg}},$ (20)
with $i$ being the unit imaginary number.
@@ II.2 Electric case
$F^{10}=\frac{q_{e}}{L_{F}r^{2}}\sqrt{\frac{f}{g}},\quad\mbox{and}\quad F=-\frac{2q_{e}^{2}}{r^{4}L_{F}^{2}},$ (21)
where $q_{e}$ is the electric charge.
@@
$-\frac{f^{\prime}}{r}-\frac{f}{r^{2}}+\frac{1}{r^{2}}=\frac{L}{2}+\frac{2q_{e}^{2}}{r^{4}L_{F}}+\epsilon f\phi^{\prime 2}+V,$ (22)
$-\frac{f\left(\frac{rg^{\prime}}{g}+1\right)}{r^{2}}+\frac{1}{r^{2}}=\frac{L}{2}+\frac{2q_{e}^{2}}{r^{4}L_{F}}-\epsilon f\phi^{\prime 2}+V,$ (23)
@@
$L_{F}(r)=\frac{8q_{e}^{2}g^{2}}{r^{2}\left(rg\left(rf^{\prime}g^{\prime}-2f\left(g^{\prime}-rg^{\prime\prime}\right)\right)+2g^{2}\left(rf^{\prime}+f\left(-4r^{2}\phi^{\prime 2}-2\right)+2\right)-r^{2}fg^{\prime 2}\right)},$ (26)
@@
Although obtained independently through the Einstein equations, the electromagnetic functions $L$ and $L_{F}$, for both the electric and magnetic cases, are related to each other, since $L_{F}$ is the derivative of the electromagnetic Lagrangian with respect to the electromagnetic invariant, and must obey the consistency relation:
$L_{F}=\frac{\partial L}{\partial F}=\frac{\partial L}{\partial r}\left(\frac{\partial F}{\partial r}\right)^{-1}\rightarrow L_{F}-\frac{\partial L}{\partial r}\left(\frac{\partial F}{\partial r}\right)^{-1}=0.$ (27)
@@
$P=P^{\mu\nu}P_{\mu\nu}=L_{F}^{2}F=-\frac{2q_{e}^{2}}{r^{4}}.$ (28)
In general, the above function is much easier to invert than the scalar $F$, the equation (21).
@@ III Simpson–Visser Solution
In order to test the consistency of the method, we will first consider the Simpson-Visser BB solution, which has a bounce structure similar to LQG-inspired solutions. Usually, the line element that describes this spacetime is written as [Simpson:2018tsi]
@@
Despite presenting itself differently from the equation (1), through a coordinate transformation, we can rewrite this line element as the equation (1) with the functions $f(r)$ and $g(r)$ given by [Canate:2022gpy]
@@
$L(r)$ $=$ $\frac{12a^{2}m}{5r^{5}},\quad L_{F}(r)=\frac{3a^{2}m}{2q_{m}^{2}r},\quad F(r)=\frac{2q_{m}^{2}}{r^{4}}.$ (33)
With appropriate adaptations of the coordinates and notation, the results are compatible with those obtained in previous works [Canate:2022gpy, Bronnikov:2021uta, Rodrigues:2023vtm]. The electromagnetic quantities satisfy the relation (27).
@@
$L(F)=\frac{3\ 2^{3/4}a^{2}F^{5/4}m}{5q_{m}^{5/2}},\qquad V(\phi)=\frac{4m}{5a^{3}\sec(\phi)^{5}}.$ (34)
Expanding $V(\phi)$ to $\phi\approx 0$, we find
$V(\phi)\approx\frac{4m}{5a^{3}}-\frac{2m\phi^{2}}{a^{3}}.$ (35)
Thus, we see that at zero the potential tends to a constant, and the first correction term is proportional to $\phi^{2}$, similar to the correction of a massive scalar field potential. In Fig. 1, we see the behavior of the potential, which resembles a potential barrier that grows as the value of $a$ decreases.
Figure 1: Behavior o the scalar field potential to the Simpson-Visser solution as a function of $\phi$ considering different values of $a$.
@@
$L(r)=-\frac{18a^{2}m}{5r^{5}},\quad L_{F}(r)=\frac{2q_{e}^{2}r}{3a^{2}m},\quad F(r)=-\frac{9a^{4}m^{2}}{2q_{e}^{2}r^{6}}.$ (36)
@@
Once again, the result is consistent with those already existing in the literature [Alencar:2024yvh]. This shows us that the method is effective in obtaining the source of some types of BBs in GR.
It is important to emphasize that in some cases it is not possible to represent the source of solutions using a scalar field that is purely canonical or purely phantom, thus requiring the consideration of partially phantom scalar fields [Bronnikov:2022bud, Bolokhov:2024sdy, Crispim:2024lzf, Rodrigues:2025plw, GRAVMOD4]. This type of model demands a greater degree of freedom in the scalar field, thereby introducing additional functions to be determined.
@@ IV Holonomy corrected Schwarzschild BH
The Holonomy corrected Schwarzschild BH is described by the line element (1) with [Alonso-Bardaji:2021yls]
@@
where $0<r_{0}<2m$ represents the radius of the throat of a WH with minimal area $4\pi r_{0}^{2}$ [Moreira:2023cxy].
Despite the metric coefficient being different, the causal structure of the Simpson-Visser solution is very similar to Holonomy corrected Schwarzschild BH. Therefore, it is possible that this spacetime, like the Simpson-Visser solution, can also be found from GR with the appropriate sources. Motivatingly, the geometry of this spacetime was investigated with regard to gravitational lensing effects, as well as the deflection of light into the strong and weak field regimes [Soares1]. Subsequently, these analyses were extended to a system where the geometry is topologically charged due to the presence of a global monopole [Soares2, Ahmed1]. Gravitational effects of this geometry have also been explored in the study of source field stability as well as usual and long-lived quasinormal modes [MODOS1, MODOS2], as well as in the context of tidal effects for a regularized geometric version [Crispim:2025cql].
@@
$L(r)$ $=$ $\frac{r_{0}(9m+4r)}{6r^{4}},\quad L_{F}(r)=\frac{r_{0}(3m+r)}{4q_{m}^{2}},\quad F(r)=\frac{2q_{m}^{2}}{r^{4}}.$ (40)
@@
$L(F)=\frac{r_{0}\left(4\sqrt[4]{2}F^{3/4}\sqrt{q_{m}}+9Fm\right)}{12q_{m}^{2}},\qquad V(\phi)=\frac{\cos^{6}\left(\frac{\phi}{\sqrt{2}}\right)\left(3m\cos\left(\sqrt{2}\phi\right)+3m+4r_{0}\right)}{24r_{0}^{3}}.$ (41)
In the limit of weak fields, $F\approx 0$, the dominant term is $F^{3/4}$, which is nonlinear. Expanding $V(\phi)$ to $\phi\approx 0$, we find
$V(\phi)\approx\frac{6m+4r_{0}}{24r_{0}^{3}}-\frac{\phi^{2}(2m+r_{0})}{4r_{0}^{3}}.$ (42)
Thus, we see that at zero the potential tends to a constant, and the first correction term is proportional to $\phi^{2}$, similar to the correction of a massive scalar field potential, similar to the Simpson-Visser case but with a different mass. In Fig. 2, we see the behavior of the potential, which resembles a potential barrier that grows as the value of $r_{0}$ decreases.
Figure 2: Behavior o the scalar field potential as a function of $\phi$ considering different values of $r_{0}$.
@@
$L(r)=-\frac{r_{0}(9m+2r)}{6r^{4}},\quad L_{F}(r)=\frac{4q_{e}^{2}}{3mr_{0}+rr_{0}},\quad F(r)=-\frac{r_{0}^{2}(3m+r)^{2}}{8q_{e}^{2}r^{4}}.$ (43)
@@
$L(P)=\frac{r_{0}\left(9mP-2\sqrt[4]{2}(-P)^{3/4}\sqrt{q_{e}}\right)}{12q_{e}^{2}}.$ (44)